OpenStudy (realitywillslapyou):

Please help me! I am running out of time and I need to have this done! Consider the line that passes through the point (3, -2) and has a slope of 2. Write the equation of this line using point-slope form. Thank you so much!!

OpenStudy (anonymous):

So point slope form is y – y1 = m(x – x1) where m is the slope and (x,y) are x and y -2 – y1 = 2(3 – x1)

OpenStudy (realitywillslapyou):

Thank you so much!! :3 ~ <3

OpenStudy (realitywillslapyou):

Can you help me just a little more?

OpenStudy (realitywillslapyou):

Using the equation from the question above, rewrite this equation in slope-intercept form. Make sure to show all of your work.

OpenStudy (anonymous):

So slope intercept form is y=mx+b Where m is the slope, (x,y) are x and y, and b is this random constant. If you substitute the slope in, you get y=2x+b Substituting the x and y gets you -2=2(3)+b Solve the equation for the b b=-8 then plug-in the slope and b into the original slope intercept form y=2x-8

OpenStudy (realitywillslapyou):

Thank you!! You are so good!

OpenStudy (realitywillslapyou):

Can you hep me with just one more? Please!

OpenStudy (anonymous):

Hurry.

OpenStudy (realitywillslapyou):

Using the equation from just above, rewrite this equation in standard form. Make sure to show all of your work

OpenStudy (anonymous):

So standard form is basically just shoving the x and y to one side of the equal sign and a constant on the other. From y=2x-8, you subtract the 2x to the other side to get y-2x=-8. If you want to clean-up, 2x-y=8

OpenStudy (anonymous):

Note that I'm not sure about standard form because it is seldom used after integrated algebra.

OpenStudy (realitywillslapyou):

That's okay! Don't worry!

OpenStudy (realitywillslapyou):

Thank you soooooo very very much!!! You have saved me!! *yay!* I would give you right now a hug if I could! :D

OpenStudy (anonymous):

mhmm, nty.

OpenStudy (realitywillslapyou):

Good Night!